OpenStudy (anonymous):

I Am Awarding Medals! @satellite73 calling @amistre64 @Hero, UNITE! What is the simplified form of the expression 8the square root of v squared. − the square root of v to the fourth power. − the square root of the quantity 9 times v squared. + the square root of the quantity 16 times v to the fourth power.

OpenStudy (anonymous):

The answer choices are 15v2 + 11v 11v2 + 15v 5v2 + 3v 3v2 + 5v

OpenStudy (anonymous):

@Hero , if you're not too busy do you think you could help?

OpenStudy (anonymous):

Or @amistre64 you're good too.

OpenStudy (amistre64):

i dont get good grades in my english classes .... even tho i was borned over here

OpenStudy (anonymous):

What other language do you speak

OpenStudy (amistre64):

i dabble in a few; even samerican sign language, but im not proficient at any of them. English is my strongest tho

OpenStudy (anonymous):

Ok, either you'll be fine, I'm sure I'll be able to understand

OpenStudy (amistre64):

8the square root of v squared: \(8\sqrt{v^2}\) − the square root of v to the fourth power: \(-(\sqrt v)^4\) − the square root of the quantity 9 times v squared: \(-\sqrt{9v^2}\) + the square root of the quantity 16 times v to the fourth power: \(+\sqrt{16v^4}\)

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

\[\Large 8v^{2/2}-v^{4/2}-3v^{2/2}+4v^{4/2}\]

OpenStudy (anonymous):

That's not the equation

OpenStudy (amistre64):

oh, then i guess someone else needs to do it ....

OpenStudy (anonymous):

The equation look like this

OpenStudy (amistre64):

@Hero any ideas?

hero (hero):

Maybe @santiagoto should write what he thinks the equation is.

OpenStudy (anonymous):

\[8\sqrt{v^2}-\sqrt{v^4}-\sqrt{9v^2}+\sqrt{16v^4}\]

OpenStudy (amistre64):

a few propeties to keep in mind:\[\sqrt{ab}=\sqrt{a}~\sqrt{b}\\~\\\sqrt{k^n}=k^{n/2}\]

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

\[8v^{2/2}-v^{4/2}-3v^{2/2}+4v^{4/2}\]

OpenStudy (anonymous):

@Hero can you show me on vyew, that really helped

OpenStudy (anonymous):

Hello?

hero (hero):

Try factoring out common factors.

hero (hero):

What's common to each term?

OpenStudy (anonymous):

\[v ^{2}\]?

hero (hero):

\[\sqrt{v^2}\]

OpenStudy (anonymous):

Oh

hero (hero):

But that simplifies to something

OpenStudy (anonymous):

So \[8-\sqrt{v^2}-3+\sqrt{16v^2}\]

hero (hero):

What happened to the expression you factored out?

OpenStudy (anonymous):

I divided everything by sqrt{v^2}

hero (hero):

You have to place it on the outside with parentheses around the expression above.

hero (hero):

\[\left(8 - \sqrt{v^2} - 3 \sqrt{16v^2}\right)\sqrt{v^2}\]

hero (hero):

What does sqrt(v^2) simplify to?

OpenStudy (anonymous):

I was typing that equation in. I thought that Sqrt{v^2} goes in the front

hero (hero):

It doesn't matter whether you put in front or back, as long as it's factored out bro. now what does sqrt(v^2) simplify to? C'mon bro.

OpenStudy (anonymous):

v

OpenStudy (anonymous):

??

hero (hero):

Yes. Simplify all of the squares.

hero (hero):

Remember the rule \[\sqrt{ab} = \sqrt{a} \dot\ \sqrt{b}\]

OpenStudy (anonymous):

Wait, that's kinda confusing

hero (hero):

\[\sqrt{16v^2} = \sqrt{16} \dot\ \sqrt{v^2}\] What's confusing about that?

hero (hero):

It's the same exact thing I showed you yesterday

OpenStudy (anonymous):

Well it was confusing to find the difference of a and b you never showed me that, or did you?

hero (hero):

I showed you that rule yesterday.

OpenStudy (anonymous):

Oh yeah, lemme get my notes

OpenStudy (anonymous):

Got em

hero (hero):

Anyway, it simplifies to \[(8 - v - 12v)v\] You should be able to finish simplifying it from there.

OpenStudy (anonymous):

16

OpenStudy (anonymous):

!!

OpenStudy (anonymous):

Is that right? @Hero

hero (hero):

I don't know where you get 16 from bro. All you had to do is simplify the expression. It doesn't simplify to a single number.

hero (hero):

The expression will include varable v, but it will be in a more simplified form.

OpenStudy (anonymous):

I kept going to try other questions and I still don't get it helppppp! @Hero

hero (hero):

I thought you were understanding this but I was wrong. If you can't simplify the expression above then you need more help than previously thought.

OpenStudy (anonymous):

Not really maybe a quick review anything you got. I understand some but then I get lost and I don't see how some things are done that's when I get tripped up

hero (hero):

I don't see why you cannot simplify (8 - v - 12v)v

hero (hero):

There's nothing to get tripped up about it. That's basic algebra.

OpenStudy (anonymous):

I can do that! Just I was confused before. Remember you're already used to this I'm still just learning it. I'm a rookie and you're a college math guru.

hero (hero):

If you can do it, then what is the simplifed version of that expression?

hero (hero):

You wrote "16" before which means nothing in terms of simplifying the expression.

OpenStudy (anonymous):

8v-v^2-12v^2

hero (hero):

And you cannot simplify it further?

OpenStudy (anonymous):

8v+12v^4?

hero (hero):

No. there's nothing about the expression that suggests to multiply anything.

OpenStudy (anonymous):

Isnt that why the v is in the front, for multiplication

hero (hero):

7 - 4 = 3 7x - 2x = 5x 5x^2 - 2x^2 = 3x^2

hero (hero):

See a pattern there?

hero (hero):

When you have like terms you combine them.

OpenStudy (anonymous):

No :?

hero (hero):

Well, I do...and it's obvious.

hero (hero):

But here's the more interesting thing.

hero (hero):

7x - 2x = (7 - 2)x = 5x 5x^2 - 2x^2 = (5 - 2)x^2 = 3x^2

OpenStudy (anonymous):

Oh

hero (hero):

Remember when we were talking about factoring out what's common?

hero (hero):

(8 - v - 12v)v has two terms with a common variable. Combine them together first: (8 - 13v)v You can leave it like that or you can multiply the v to get 8v - 13v^2

hero (hero):

Either of those last two expressions are acceptable for a final answer.

OpenStudy (anonymous):

Oh, I get it thanks again. And you work at a college?

OpenStudy (anonymous):

And unfortunately, that isn't one of the answer choices, now what?!

hero (hero):

It's not one of the answer choices because I made a mistake.

hero (hero):

It's hard not to make a mistake while typing on this computer.

OpenStudy (anonymous):

So that's why I got tripped up lol, I knew something was wrong.

hero (hero):

My mistakes are usually minor mistakes that could lead to huge mistakes. But your mistakes are just based on not knowing.

OpenStudy (anonymous):

Whats up with your computer

hero (hero):

There's nothing up with it. I just type too fast sometimes and miss a plus or minus here and there.

hero (hero):

Which is not the same as you not understanding how to simplify an expression.

hero (hero):

I guess now we can test that out to see if you really understand.

OpenStudy (anonymous):

Ok lets go for it. :/

hero (hero):

I can show you the mistake I made. I left out a plus: Incorrect: \[\left(8 - \sqrt{v^2} - 3 \sqrt{16v^2}\right)\sqrt{v^2}\] correct: \[\left(8 - \sqrt{v^2} - 3 + \sqrt{16v^2}\right)\sqrt{v^2}\] Now you can try simplifying the correct expression.

OpenStudy (anonymous):

(5-sqrt{v^2}+sqrt{16v^2})sqrt{v^2} So far

hero (hero):

Bro Did you already forget what sqrt(v^2) reduces to?

hero (hero):

And you also seemed to have forgotten what sqrt(ab) equals

OpenStudy (anonymous):

v, I know I'm taking it step by step bro

hero (hero):

Well, post each step line by line all in the same post. Then submit.

OpenStudy (anonymous):

ok (8-v-3+sqrt{16v})sqrt{v} (5+sqrt{17v})sqrt{v}

hero (hero):

That's not correct.

hero (hero):

\[\sqrt{v^2} \ne \sqrt{v}\]

hero (hero):

Also \[\sqrt{16v^2} \ne \sqrt{16v}\]